what are the factors of the number 2400? and how many are there in total?
Answer by Edwin McCravy(20055)

what are the factors of the number 2400? and how many are there in total?


Break 2400 down into prime factors with a factor tree:

2400 = 2·2·2·2·2·3·5·5

It's easy to answer the last question.  There are only
3 prime factors of 2400. These are 2,3,and 5.  So
every factor is of the form

2P3Q5R

where P can take on values 0,1,2,3,4,5, that's 6 factors
      Q can take on values 0,1, that's 2 factors
      R can take on values 0,1,2, that's 3 factors

So that's a total of 6·2·3 = 36 factors.

But to get all those 36 factors is a lot of busy work:

We make all the combinations of those factors:

--------------------------------------

Combinations of one factor:

1, 2, 3, and 5

This accounts for 4 factors

-----------------------------

Combinations of two factors:

2·2, 2·3, 2·5, 3·5. 5·5 or

This accounts for 5 factors

-----------------------------

Combination of three factors:

2·2·2, 2·2·3, 2·2·5, 2·3·5, 2·5·5, 3·5·5

This accounts for 6 factors

----------------------------

Combinations of four factors:

2·2·2·2, 2·2·2·3, 2·2·2·5, 2·2·3·5, 2·2·5·5, 2·3·5·5 

This accounts for 6 factors

-----------------------------

Combinations of five factors:

2·2·2·2·2, 2·2·2·2·3, 2·2·2·2·5, 2·2·2·3·3, 2·2·2·3·5, 2·2·2·5·5, 2·2·3·5·5

This accounts for 7 factors

------------------------------

Combinations of six factors:

2·2·2·2·2·3, 2·2·2·2·2·5, 2·2·2·2·3·5, 2·2·2·3·5·5

This accounts for 4 factors

------------------------------

Combinations of seven factors:

2·2·2·2·2·3·5, 2·2·2·2·2·5·5, 2·2·2·2·3·5·5

This accounts for 3 factors

-------------------------------

Combinations of eight factors:

2·2·2·2·2·3·5·5

This accounts for 1 factor.

--------------------------------

That's  4 + 5 + 6 + 6 + 7 + 4 + 3 + 1 = 36 factors, so we have them
all. These are, in order:

1, 2, 3, 4, 5, 6, 8, 10, 12, 15, 16, 20, 24, 25, 30, 32, 40, 48, 50,
60, 75, 80, 96, 100, 120, 150, 160, 200, 240, 300, 400, 480, 600, 800,
1200, 2400

Edwin